Two people are jogging around a circular track in the same direction. One person can run comepletely around the track in 15 minutes. The second person takes 18 minutes. If they both start running in the same place at the same time, how long will it take them to be together at this place if they continue to run?
Found 2 solutions by sudhanshu_kmr, stanbon:
Answer by sudhanshu_kmr(1152)


Find the LCM of 15 and 18.
it is 90
so, after 90 minutes they will together at the place.

Answer by stanbon(75887)

Two people are jogging around a circular track in the same direction. One person can run comepletely around the track in 15 minutes. The second person takes 18 minutes. If they both start running in the same place at the same time, how long will it take them to be together at this place if they continue to run?
----
Find the least common multiple of 15 and 18:
15 = 3*5
18 = 3^2*2
------
lcm = 3^2*2*5 = 90
-----
It will take 90 minutes for them to be together.
Cheers,
Stan H.